The area of a circle is equal to the area of a square measuring 5 centimeters on each side. Find the radius if the circle.

Answer by harpazo(655)

From the information given, the area of the square turns out to be 25.
This is the same area of the circle.
25 = pi*r^2
Solve for r.
25/pi = r^2
Now take square root of both sides.

sqrt{25/pi} = sqrt{r^2}
5/sqrt{pi} = r

Did you follow?